Problem 22-03
Merger Bid

Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.50 (given its target capital structure). Vandell has $10.04 million in debt that trades at par and pays an 7.5% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 6% a year. Both Vandell and Hastings pay a 35% combined federal and state tax rate. The risk-free rate of interest is 6% and the market risk premium is 6%.

Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.6 million, $2.9 million, $3.3 million, and $3.95 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 6% rate. Hastings plans to assume Vandell’s $10.04 million in debt (which has an 7.5% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.5 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.460 million, after which the interest and the tax shield will grow at 6%.

Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations.

The bid for each share should range between $_____ per share and $_____ per share.
Answer #1

Ans.

Calculation of WACC to discount Vandell's future cash flows:

Cost of Equity as per Capital Asset Pricing Model ( CAPM )

Ke = Rf + ( Beta* Market Risk premium)

Ke = 6% + ( 1.5 * 6%) = 6% + 9% = 15%

After-tax cost of debt = 7.5%*(1- 35%) = 7.5% * 0.65 = 4.875%

So,

WACC = (We*Ke) + (Wd*Kd)

WACC = (70%*15%)+(30%*4.875%)

WACC = 10.5% + 1.4625% = 11.9625% or 11.96%

Before synergy

Value of Vandell using Constant-growth FCF model by discounting at the above WACC (11.96%) & finding its PV of future cash flows

FCF1/(r-g) = (FCF0*(1+g))/(r-g)

Where,

g= the constant growth rate = 6%

r= the required return or WACC =11.96%

FCF0 = $2,000,000

By putting the values,

$2,000,000*1.06/(11.96% - 6%) = $ 2,120,000 / 5.96% = $35,570,469.79

So,Firm Value of Vandell before considering synergy= $35,570,469.79

Value of Debt = $ 10,400,000

So, value of Equity=  $35,570,469.79 - $ 10,400,000 = $ 25,170,469.79

No.of equity shares o/s =1,000,000

So, Value /share= $ 25,170,469.79/1,000,000 = 25.17

After synergy:

Year 0 1 2 3 4
Free Cash Flow (FCF) 2.6 2.9 3.3 3.95
Terminal FCF(3.95*1.06)/(11.96%-6%) 70.2517
Less: Interest payments -1.5 -1.5 -1.5 -1.46
Less: Terminal Int.exp.(1.460*1.06)/(11.96%-6%) -25.9664
Add: Interest tax shields(Int.exp.*Tax rate) 0.525 0.525 0.525 0.511
Add: Terminal Interest Tax Shield 9.08824
FCF to Equity/FCFE 1.625 1.925 2.325 56.37454
PV F at 11.96%(1/1.1196^n) 0.893176 0.797764 0.712543 0.636427
PV at 11.96% 1.451411 1.535695 1.656663 35.87827
NPV or Value to Equity 40.522036
Value of Vandell's Equity after synergy = $40,522,036
No.of equity share o/s 1000000
Value/share $ 40.52

So,

The bid for each share should range between $ 25.17 per share and $ 40.52 per share.